### v1

@@ 1 Introduction
Figure 1: Overview of the centralized seed fulfillment process. The process begins with the arrival of seed stocks from multiple sites with stochastic, a priori unknown arrival distributions and ends with the fulfillment of orders with different deadlines and quantities. Our proposed adaptive adaptive hybrid tree search approach provides an efficient solution to the wave scheduling problem, optimizing the process of order fulfillment.
Order fulfillment in industries such as e-commerce [2] and retail [3] often involve centralized fulfillment centers that simultaneously process arriving inventory and fulfill orders based on their deadlines. The fulfillment process with large catalogs often handle a batch of orders, hereinafter referred to as wave, together using automated sortation systems [4]. The supply chain in these sectors is typically well-established, with known inventory quantities and deterministic restock times. The problem of optimally scheduling waves to maximize fulfillment efficiency is addressed using traditional operations research and optimization techniques [5], [6] as order deadlines and inventory levels are known a priori or can be forecasted with low uncertainty.
However, the seed distribution system presents a unique set of challenges that are not typically encountered in traditional supply chains. The nature of seed development and distribution involves a plethora of seed varieties, each with specific growing conditions and market demands. The perishability and seasonality of seeds make seed supply chains a highly intricate web of moving parts [7], [8]. Complicating matters further, the seed supply chain is inherently stochastic – the variability in seed yields is influenced by numerous factors including weather conditions, soil quality, environmental factors [9], crop diseases [10], [11], and pest infestations [12], [13].
@@
Markov Decision Processes (MDPs) [17], [18] serve as a useful framework to approach stochastic decision-making problems. MDPs have been widely applied in a range of domains such as robotics [19], [20], [21], finance [22], [23], and healthcare [24], providing a systematic way to make optimal decisions in situations characterized by uncertainty. However, directly applying traditional solution methods such as value iteration [25] and policy iteration [26] to wave scheduling problem with large state and action spaces is computationally intractable due to the curse of dimensionality [27]. Instead, Monte Carlo tree search (MCTS) [28] provides an intelligent tree search method that is effective in large MDPs where exhaustive search is not feasible [29]. However, the size of the search space in the problem of seed order fulfillment poses challenges even for MCTS.
To address this challenge, we propose an adaptive hybrid tree search approach that augments MCTS with domain-specific knowledge to efficiently navigate the decision space. The key insight is that we can augment MCTS with additional problem-specific side information to make it tractable for large MDPs. The proposed hybrid tree search method dynamically narrows down the action space by leveraging the information about order deadlines and inventory levels. Our approach is particularly suited for scenarios with large state and action spaces, as it iteratively constructs a practical search tree that helps in exploring the most promising strategies.
@@ 2.3 Markov Decision Processes
$V_{H}^{\pi}(s)=\mathbb{E}\left[\sum_{k=0}^{H}\mathcal{R}(s_{k},\pi(s_{k}),s_{k+1})\bigg{|}s_{0}=s\right].$
@@
$Q_{H}^{\pi}(s,a)=\mathbb{E}\left[\sum_{k=0}^{H}\mathcal{R}(s_{k},\pi(s_{k}),s_{k+1})\bigg{|}s_{0}=s,\pi(s_{0})=a\right].$
@@ 3 Problem Formulation
Formally, let $T$ represent the total number of discrete time steps in the planting season, each with a duration of $\Delta t$. Consider a set of products $P=\{p_{1},p_{2},\ldots,p_{n_{p}}\}$ in the catalog. We assume that the quantity of each product replenished in the inventory at any time step is stochastic and unknown a priori. As a consequence, the total incoming quantity of all products is stochastic. Let $q^{p}_{t}\in\mathbb{Z}$ denote the quantity of product $p$ that arrived at time step $t$.
Figure 2: Illustration of the order fulfillment process in the warehouse with (a) incoming product quantities, (b) items placed in intermediate containers, (c) orders fulfilled by the system.
@@
In order to handle situations where it is not feasible to satisfy all order deadlines, we introduce a pre-emptive processing parameter, $\alpha$, ranging from $0.5$ to $1$. The parameter $\alpha$ indicates the required proportion of available items in an order relative to the total number of items for initiating processing. Importantly, the preemptive processing is reserved for the last feasible opportunity to process an order without missing its deadline. For instance, if $\alpha$ is $0.5$, order processing is triggered only when at least half of the items are available and processing at this point is the only way to avoid missing the deadline. This approach ensures the timely, albeit partial, fulfillment of orders, aligning closely with their respective deadlines.
@@
{mdframed}
[linewidth=1pt] Problem Formulation (Wave Scheduling Problem)
Given a set of orders and incoming inventory in a warehouse setting, minimize the total delay across all orders, defined as
$\min_{\mathbf{w}}\ \ D=\sum_{i=1}^{n_{o}}\max(0,f_{i}(\mathbf{w},\alpha)-d_{i})$ (1)
while satisfying the constraints ensuring that the required inventory is available to fulfill the orders, complete and preempted, at the beginning of each wave, each order is assigned to exactly one wave, and the number of orders in a wave is less than or equal to the wave capacity.
@@ 4.1 Stochastic Modeling of Product Arrival Distributions
Let $E_{p,t}^{j}(s\rightarrow s^{\prime})$ represent the transition from state $s$ at time $t-1$ to state $s^{\prime}$ at time $t$ during the $j$-th season for product $p$. Similarly, let $E_{p,t}^{j}(s)$ be the event that the state at time $t-1$ was $s$. The transition probability is then given by $\mathcal{P}_{p,t}(s,s^{\prime})=\frac{\sum_{j=1}^{m_{p}}\#(E_{p,t}^{j}(s\rightarrow s^{\prime}))}{\sum_{j=1}^{m_{p}}\#(E_{p,t}^{j}(s))}$ where #$\#(E_{p,t}^{j}(s\rightarrow s^{\prime}))$ denotes the number of times the event $E_{p,t}^{j}(s\rightarrow s^{\prime})$ occurred and #$\#(E_{p,t}^{j}(s))$ denotes the number of times the event $E_{p,t}^{j}(s)$ occurred.
@@
To illustrate the model’s effectiveness in modeling crop harvest dynamics, we employ the historical data on the harvest progress of corn in Iowa obtained from the United States Department of Agriculture (USDA) [37], as shown in Figure 3(a). We use the data from 2015 to 2022 to estimate the transition probabilities for the time-varying Markov chain model. Figure 3(b) shows ten simulated instances of harvest progression, demonstrating the model’s effectiveness in accurately modeling these agricultural trends.
@@ 4.2 Wave Scheduling Problem as a Markov Decision Process
$\mathcal{R}(s,a,s^{\prime})=\sum_{o_{i}\in O_{a}}\mathbb{I}[f_{i}\leq d_{i}]-\lambda\sum_{o_{i}\in O_{a}}\max(0,f_{i}-d_{i})$
@@ 4.3 Adaptive Hybrid Tree Search for Wave Scheduling
At any decision step in an episode, let $O_{uf}$ denote the set of unfulfilled but feasible orders, i.e., orders that have not been fulfilled but can be fulfilled given the current inventory levels. In order to reduce the action space size, we generate two subsets $O_{uf}^{d}$ and $O_{uf}^{p}$ from $O_{uf}$. We select orders for $O_{uf}^{d}$ based on the ascending deadline order while ensuring that orders past their deadlines do not account for more than a predetermined proportion of the total number of orders in $O_{uf}^{d}$. For $O_{uf}^{p}$, we first calculate the peak of unfulfilled order deadlines, i.e., the time step at which the maximum number of orders are unfulfilled, and select orders based on their proximity to this peak. We constrain $|O_{uf}^{d}|\leq 2n_{w}$, $|O_{uf}^{p}|\leq 2n_{w}$, and $|O_{uf}^{d}\cap O_{uf}^{p}|=\emptyset$. The intuition is that $O_{uf}^{d}$ prioritizes orders with earlier deadlines, while $O_{uf}^{p}$ focuses on orders likely to cause bottlenecks due to deadlines clustering around the peak. This decomposition attempts to balance urgency of individual orders against long-term risk of developing system-wide bottlenecks.
To construct the candidate order set $O_{c}$ of size $2n_{w}$, we introduce a dynamic parameter called the peak reducing factor $\rho\in[0,1]$. The value of $\rho$ varies based on tree search evaluation outcomes from previous rounds; details on how $\rho$ adapts are provided in later sections. $\rho$ adaptively adjusts the proportion of orders selected from $O_{uf}^{d}$ and $O_{uf}^{p}$ while ensuring a minimum number from each set. $O_{c}$ consists of the first $2n_{w}\rho$ orders from $O_{uf}^{d}$ and the first $2n_{w}(1-\rho)$ orders from $O_{uf}^{p}$. When $|O_{uf}^{d}|<2n_{w}\rho$, we add the remaining orders from $O_{uf}^{p}$ to $O_{c}$, and conversely, when $|O_{uf}^{p}|<2n_{w}(1-\rho)$, we add the remaining orders from $O_{uf}^{d}$ to $O_{c}$.
@@
$\pi_{tree}(s)=\operatorname*{arg\,max}_{a\in A}\left[\hat{Q}(s,a)+c\sqrt{\frac{\log\#(s)}{\#(s,a)}}\right],$
where $\hat{Q}(s,a)$ is the current estimate of state-action value, #$\#(s)$ is the visit count of state $s$, #$\#(s,a)$ is the visit count of state-action pair $(s,a)$, and $c$ is a constant that controls the amount of exploration.
@@
3:Initialize state-action rewards $Q$ and counts ##\##
@@
12:     Select actions from $\mathcal{A}_{s}$ using $\pi_{tree}$ until leaf code
@@
17:     Update $Q$ and ##\## for along the path from root to leaf node
@@ 5 Experiments and Results
In this section, we present the results of our experiments comparing the baseline ‘greedy’ approach and the proposed approach. We begin with a description of our experimental setup, baseline policy, and then proceed to discuss the results.
@@ 5.1 Experimental Setup
We conduct our experiments in a simulation environment designed to replicate the operational dynamics of a real-world centralized seed processing facility. While the simulated facility information is not publicly accessible, we expect the results to be consistent on any simulator that can model the problem. The simulator accurately models the core components of the warehouse operations, including inventory management, order processing, and the automated sortation system. The environment is parameterized to align closely with real-world data, ensuring the relevance and applicability of our findings.
We consider a season duration of 90 days which is consistent with typical seed distribution season lengths. The time step size is set to 1 minute to capture the fine-grained information in the sortation system. As discussed in Section 3, the sortation system receives the inventory in intermediate containers. We consider intermediate containers with a capacity of 250 items and assume that the facility has sufficient capacity to accommodate all the intermediate containers. The sortation system has 30 induction stations and 400 chutes, which restricts the maximum number of orders processed in a wave also to 400. An order is fulfilled when all the items in the order reach a designated chute and a wave ends only when all of its orders are fulfilled.
The experiment simulates 200 products in the catalog, each with a unique arrival distribution over the season. The quantity of each product arriving over the season is consistent with the quantity needed to fulfill the all the orders. We model the arrival distribution of the products over the season as a bimodal Gaussian distribution with the first peak around the beginning of the season and the second peak around the middle of the season. This distribution reflects the real-world scenario where the season begins with incoming product inventory from the previous season stored at several location and the new inventory from current harvest starts arriving around slowly with a peak around the middle of the season. Figure 5 illustrates the arrival distribution of two different products over the season. We assume that the inventory arrives at every 8 hours, which is consistent with the real-world setting.
@@
5.2 Baseline Approach

### v2

@@ Abstract
journal: IFAC EAAIlabel1label1footnotetext: This work was supported by Corteva Agriscience under the research collaboration agreement between Corteva Agriscience and the University of Illinois Urbana-Champaign.
@@ 1 Introduction
Figure 1: Overview of the centralized seed fulfillment process. The process begins with the arrival of seed stocks from multiple sites with stochastic, a priori unknown arrival distributions and ends with the fulfillment of orders with different deadlines and quantities. Our proposed adaptive hybrid tree search approach provides an efficient solution to the wave scheduling problem, optimizing the process of order fulfillment.
Order fulfillment in industries such as e-commerce [2] and retail [3] often involve centralized fulfillment centers that simultaneously process arriving inventory and fulfill orders based on their deadlines. The fulfillment process with large catalogs often handles a batch of orders, hereinafter referred to as wave, together using automated sortation systems [4]. The supply chain in these sectors is typically well-established, with known inventory quantities and deterministic restock times. The problem of optimally scheduling waves to maximize fulfillment efficiency is addressed using traditional operations research and optimization techniques [5], [6] as order deadlines and inventory levels are known a priori or can be forecasted with low uncertainty.
However, the seed distribution system presents a unique set of challenges that are not typically encountered in traditional supply chains. The nature of seed development and distribution involves a plethora of seed varieties, each with specific growing conditions and market demands. The perishability and seasonality of seeds make seed supply chains a highly intricate web of moving parts [7], [8]. Recent advances in agricultural supply chain management [9, 10] have highlighted the growing complexity of seed distribution networks. Complicating matters further, the seed supply chain is inherently stochastic – the variability in seed yields is influenced by numerous factors including weather conditions, soil quality, environmental factors [11], crop diseases [12], [13], and pest infestations [14], [15].
@@
Markov Decision Processes (MDPs) [19], [20] serve as a useful framework to approach stochastic decision-making problems. MDPs have been widely applied in a range of domains such as robotics [21], [22], [23], finance [24], [25], transportation [26], [27], healthcare [28], and large-scale infrastructure management [29], providing a systematic way to make optimal decisions in situations characterized by uncertainty. Practical considerations in MDP applications often involve resource constraints [30], which in our context translates to inventory for order fulfillment. Recent work has explored deep reinforcement learning for warehouse scheduling [31, 32], but these approaches require extensive training data that is often unavailable in agricultural settings. However, directly applying traditional solution methods such as value iteration [33] and policy iteration [34] to wave scheduling problem with large state and action spaces is computationally intractable due to the curse of dimensionality [35]. Instead, Monte Carlo tree search (MCTS) [36] provides an intelligent tree search method that is effective in large MDPs where exhaustive search is not feasible [37]. Recent applications of MCTS in logistics [38] demonstrate its effectiveness, but the exponential action spaces in wave scheduling remain challenging even for advanced MCTS variants.
To address this challenge, we propose an adaptive hybrid tree search approach that augments MCTS with domain-specific knowledge to efficiently navigate the decision space. The key insight is that we can augment MCTS with additional problem-specific side information to make it tractable for large MDPs. The novelty of our approach lies not merely in using domain knowledge, but specifically in how this knowledge is dynamically and adaptively integrated with the MCTS framework. This includes unique mechanisms such as the dual-criteria candidate order selection, the dynamic adjustment of the schedule based on MCTS feedback, and an adaptive action set generation process tailored for this domain. These elements go beyond generic heuristics, offering a structured and responsive way to manage the vast search space inherent in wave scheduling.
@@ 2.3 Markov Decision Processes
$V_{H}^{\pi}(s)=\mathbb{E}\left[\sum_{k=0}^{H}\mathcal{R}(s_{k},\pi(s_{k}),s_{k+1})\bigg|s_{0}=s\right].$ (1)
@@
$Q_{H}^{\pi}(s,a)=\mathbb{E}\left[\sum_{k=0}^{H}\mathcal{R}(s_{k},\pi(s_{k}),s_{k+1})\bigg|s_{0}=s,\pi(s_{0})=a\right].$ (2)
@@ 3 Problem Formulation
Formally, let $T$ represent the total number of discrete time steps in the planting season, each with a duration of $\Delta t$. Consider a set of products $P=\{p_{1},p_{2},\ldots,p_{n_{p}}\}$ in the catalog. We assume that the quantity of each product replenished in the inventory at any time step is stochastic and unknown a priori. As a consequence, the total incoming quantity of all products is stochastic. Let $q^{p}_{t}\in\mathbb{Z}$ denote the quantity of product $p$ that arrived at $t$.
Figure 2: Illustration of the order fulfillment process in the warehouse with (a) incoming product quantities, (b) items placed in containers, (c) orders fulfilled by the system.
@@
In order to handle situations where it is not feasible to satisfy all order deadlines, we introduce a pre-emptive processing parameter, $\alpha$, ranging from $0.5$ to $1$. This parameter, typically user-defined, represents a degree of operational flexibility afforded to the system. The parameter $\alpha$ indicates the required proportion of available items in an order relative to the total number of items for initiating processing. Importantly, this form of preemptive processing is reserved for the last feasible opportunity to process an order without missing its deadline, effectively influencing the definition of a ”feasible” order under such critical, time-sensitive conditions. Its role is to allow the system to process orders that are not fully complete but are very close to their deadline, where waiting for full availability would guarantee a missed deadline, thus balancing fulfillment rates with inventory availability. For instance, if $\alpha$ is $0.5$, order processing is triggered only when at least half of the items are available and processing at this point is the only way to avoid missing the deadline. This approach ensures the timely, albeit partial, fulfillment of orders, aligning closely with their respective deadlines.
@@
Problem Formulation (Wave Scheduling Problem) Given a set of orders and incoming inventory in a warehouse setting, minimize the total delay across all orders, defined as $\min_{\mathbf{w}}\ \ D=\sum_{i=1}^{n_{o}}\max(0,f_{i}(\mathbf{w},\alpha)-d_{i})$ (3) while satisfying the constraints ensuring that the required inventory is available to fulfill the orders, complete and preempted, at the beginning of each wave, each order is assigned to exactly one wave, and the number of orders in a wave is less than or equal to the wave capacity.
@@ 4.1 Stochastic Modeling of Product Arrival Distributions
Let $E_{p,t}^{j}(s\rightarrow s^{\prime})$ represent the transition from state $s$ at time t1t-1 to state $s^{\prime}$ at time $t$ during the $j$-th season for product $p$. Similarly, let $E_{p,t}^{j}(s)$ be the event that the state at time t1t-1 was $s$. The transition probability is then given by $\mathcal{P}_{p,t}(s,s^{\prime})=\frac{\sum_{j=1}^{m_{p}}\#(E_{p,t}^{j}(s\rightarrow s^{\prime}))}{\sum_{j=1}^{m_{p}}\#(E_{p,t}^{j}(s))}$ where $(E_{p,t}^{j}(s\rightarrow s^{\prime}))$ denotes the number of times the event $E_{p,t}^{j}(s\rightarrow s^{\prime})$ occurred and $(E_{p,t}^{j}(s))$ denotes the number of times the event $E_{p,t}^{j}(s)$ occurred.
@@
To illustrate the model’s effectiveness in modeling crop harvest dynamics, we employ historical data on the harvest progress of corn in Iowa obtained from the United States Department of Agriculture (USDA) [46], as shown in Figure 3(a). This raw data was first normalized to our standardized state space $\mathcal{S}=\{0,1,\ldots,1000\}$, yielding $\tilde{h}_{p,t}^{i}$ for each past season. From these normalized historical trajectories, we empirically tallied the transition counts $(E_{p,t}^{j}(s\rightarrow s^{\prime}))$ and $(E_{p,t}^{j}(s))$ across all seasons $j$ for each time step $t$. These aggregated counts directly inform the estimation of the time-varying transition probabilities $\mathcal{P}_{p,t}(s,s^{\prime})$ using the formula provided. To generate the simulated harvest progressions in Figure 3(b), each trajectory begins at state 0 (0% harvested) at $t=0$. At each subsequent time step $t$, given the current simulated cumulative state $s_{t}$, the next state $s_{t+1}$ is sampled from the discrete probability distribution defined by the row $\mathcal{P}_{p,t}(s_{t},\cdot)$. Repeating this iterative sampling process generates a complete simulated trajectory, and multiple such repetitions produce the ensemble of curves shown, demonstrating how the transition probabilities accurately capture and simulate to dynamic harvest trends.
@@ 4.2 Wave Scheduling Problem as a Markov Decision Process
$\mathcal{R}(s,a,s^{\prime})=\sum_{o_{i}\in O_{a}}\mathbb{I}[f_{i}\leq d_{i}]-\lambda\sum_{o_{i}\in O_{a}}\max(0,f_{i}-d_{i})$ (4)
@@ 4.3 Adaptive Hybrid Tree Search for Wave Scheduling
At any decision step in an episode, let $O_{uf}$ denote the set of unfulfilled but feasible orders, i.e., orders that have not been fulfilled but can be fulfilled given the current inventory levels. It is important to note that an order might be in $O_{uf}$ at some point but still exceed its deadline if it is not selected for processing by the algorithm; this can occur due to strategic long-term decisions made by the MCTS, such as prioritizing other more critical orders or waiting for more items to form a larger, more efficient wave, based on its lookahead capabilities. In order to reduce the action space size, we generate two subsets $O_{uf}^{d}$ and $O_{uf}^{p}$ from $O_{uf}$. Orders for $O_{uf}^{d}$ are selected based on their deadlines in ascending order, giving precedence to the most urgent ones, while ensuring that orders already past their deadlines do not constitute more than a predetermined proportion of $O_{uf}^{d}$. For $O_{uf}^{p}$, orders are selected based on their proximity to the peak of unfulfilled order deadlines (the time step with the maximum number of currently unfulfilled orders), aiming to alleviate potential future bottlenecks. If an order meets the criteria for both subsets, it is primarily considered for $O_{uf}^{d}$ due to deadline urgency; the constraint $|O_{uf}^{d}\cap O_{uf}^{p}|=\emptyset$ ensures distinctness, typically by removing an order from consideration for $O_{uf}^{p}$ if it’s already prioritized in $O_{uf}^{d}$. We constrain $|O_{uf}^{d}|\leq 2n_{w}$ and $|O_{uf}^{p}|\leq 2n_{w}$. The intuition is that $O_{uf}^{d}$ prioritizes orders with earlier deadlines, while $O_{uf}^{p}$ focuses on orders likely to cause bottlenecks due to deadlines clustering around the peak. This decomposition attempts to balance urgency of individual orders against long-term risk of developing system-wide bottlenecks.
To construct the candidate order set $O_{c}$ of size $2n_{w}$, we introduce a dynamic parameter called the peak reducing factor $\rho\in[0,1]$. The value of $\rho$ varies based on tree search evaluation outcomes from previous rounds; its adaptation based on MCTS feedback is part of the learning process discussed in Algorithm 1. $\rho$ adaptively adjusts the proportion of orders selected from $O_{uf}^{d}$ and $O_{uf}^{p}$ while ensuring a minimum number from each set. $O_{c}$ consists of the first $2n_{w}\rho$ orders from $O_{uf}^{d}$ and the first $2n_{w}(1-\rho)$ orders from $O_{uf}^{p}$. When $|O_{uf}^{d}|<2n_{w}\rho$, we add the remaining orders from $O_{uf}^{p}$ to $O_{c}$, and conversely, when $|O_{uf}^{p}|<2n_{w}(1-\rho)$, we add the remaining orders from $O_{uf}^{d}$ to $O_{c}$. This structured approach to forming $O_{c}$ by segmenting orders based on urgency and potential bottlenecks ($O_{uf}^{d}$ and $O_{uf}^{p}$) and then dynamically balancing their inclusion via $\rho$, is a key way domain-specific knowledge is integrated to prune the vast action space, making MCTS computationally tractable for this complex scheduling problem. Further refinement of the action space occurs during action set generation, where waves are constructed from $O_{c}$ and filtered by estimated fulfillment time.
@@
$\pi_{tree}(s)=\operatorname*{arg\,max}_{a\in A}\left[\hat{Q}(s,a)+c\sqrt{\frac{\log\#(s)}{\#(s,a)}}\right],$ (5)
where $\hat{Q}(s,a)$ is the current estimate of state-action value, #(s)\#(s) is the visit count of state $s$, #(s,a)\#(s,a) is the visit count of state-action pair $(s,a)$, and $c$ is a constant that controls the amount of exploration.
@@
3:Initialize state-action rewards $Q$ and counts #\#
@@
12:  Select actions from $\mathcal{A}_{s}$ using $\pi_{tree}$ until leaf node
@@
17:  Update $Q$ and #\# for along the path from root to leaf node
@@ 5 Experiments and Results
In this section, we present a comprehensive evaluation of our proposed approach against established baseline methods. We begin with a description of the experimental setup, followed by the baseline approaches used for comparison. We then present the performance results, conduct a sensitivity analysis on key parameters, analyze detailed scheduling performance by order category, and discuss the computational considerations of our method.
@@ 5.1 Experimental Setup
We conduct our experiments in a simulation environment designed to replicate the operational dynamics of a real-world centralized seed processing facility. The simulator accurately models the core components of warehouse operations including inventory management, order processing, and automated sortation, parameterized to align closely with real-world data to ensure relevance and applicability of our findings.
We consider a seed distribution season spanning 90 days, with a time step size of 1 minute to capture fine-grained information in the sortation system. As discussed in Section 3, the sortation system receives inventory in intermediate containers with a capacity of 250 items, and we assume the facility has sufficient capacity to accommodate all containers. The standard configuration features 30 induction stations and 400 chutes, which limits the maximum number of orders processed in a single wave to 400. An order is fulfilled when all its items reach the designated chute, and a wave concludes only when all its orders are fulfilled.
The simulation features 200 products in the catalog, each with a unique arrival distribution throughout the season. The quantity of each product arriving over the season is consistent with the quantity needed to fulfill all orders. We model the arrival distribution using a bimodal Gaussian distribution with peaks at the beginning and middle of the season, reflecting real-world scenarios where initial inventory comes from previous season’s storage, while new production from the current harvest peaks mid-season. Figure 5 illustrates the arrival distribution of two representative products. Inventory arrivals are scheduled every 8 hours, consistent with typical operational practices.
@@
5.2 Baseline Approaches
We compare our proposed method against three baseline approaches:
Greedy Heuristic